# Convergence Analysis of An Alternating Nonlinear GMRES on Linear Systems

### v1

@@ 4 Numerical results
In this section, we present some examples to confirm our theoretical findings, and further explore the convergence behavior of aNGMRES in cases where GMRES stagnates. In our numerical test, we use the MATLAB backslash operator as the solver for the least-squares problems. We note that when approximations closely approach the exact solution, the least-squares problems involved in NGMRES may become ill-conditioned. We leave further investigation of this issue for future study.
@@
To confirm Theorem 3.1, we take $m=3$ and $p=m+1=4$. We report the convergence history of aNGMRES(3,4) and GMRES(4). From Theorem 3.1, we know that the $(4j)$th iterate of the two methods are the same, which is confirmed in Figure 1.
Figure 1: Convergence history of aNGMRES(3,4) vs. GMRES(4) for Example 4.1.
@@
Figure 2: Convergence history for Example 4.1. Left: aNGMRES($\infty$,4) vs. GMRES. Right: aNGMRES($\infty$,5) vs. GMRES.
@@
Consider $\mathcal{A}x=b$ taken from [9], where $\mathcal{A}$ is block diagonal with $q$ blocks. Each block has the same structure as $A$ defined in (28). The $k$th block has dimension $\ell k$, where $\ell\in\mathbb{N}$.
Let us consider $\ell=3$ and $q=5$. Then, $\mathcal{A}\in\mathbb{R}^{45\times 45}(3+6+9+12+15=45)$. Note that $\mathcal{A}$ is invertible. We consider zero initial guess, and construct $b$ by setting to one all the entries whose index coincides with the row index of matrix $\mathcal{A}$ where a new block starts, i.e., the nonzero components of vector $b$ are at indexes of 1, 4, 10, 19, 31. $\nu(A,r_{0})=30$. In this example, full GMRES will stagnate for $\ell$ consecutive iterations.
@@
Figure 3: Convergence history for Example 4.2: aNGMES(2,3) vs. GMRES(3).
In Figure 4, we consider the performance of full GMRES and aNGMRES($\infty$,$p$) with $p=1,2,3,4$. Note that aNGMRES($\infty$,1) is NGMRES($\infty$). It can be seen that GMRES stagnates for three consecutive iterations, while aNGMRES($\infty$,1) stagnates forever, i.e., $u_{k}=u_{0}$ for all $k$. For aNGMRES($\infty$,2), we see the norm of the residual sequence forms an alternating periodic sequence with a period of 2. However, we see that $\|r_{3j}\|$ of aNGMRES($\infty$,3) is the same as $\|r_{3j}^{G}\|$ (i.e., $x_{3j}=x_{3j}^{G}$), and aNGMRES($\infty$,3) converges to the exact solution at the 30th iteration, which is true for GMRES. Moreover, aNGMRES($\infty$, 4) converges at the 40th iterations.
@@
Figure 4: Convergence history for Example 4.2: aNGMRES($\infty$,$p$) with $p=1,2,3,4$ vs. GMRES.
@@
We consider 2D Laplacian with finite difference scheme and construct the coefficient matrix by using MATLAB command $A=delsq(numgrid(^{\prime}S^{\prime},n+2))$ with $n=64$. We take $Au=b$ with $b$ all ones and random initial guess.
@@
In this example, the fixed-point iteration $q(u_{k+1})=Mu_{k}+b$ with $M=I-A$ does not converge. To validate Theorem 3.2, we compare aNGMRES($\infty$,$3$) with GMRES in Figure 5. We see that the subsequence $\{u_{3j}\}$ of aNGMRES($\infty$,$3$) method coincides with the sequence $\{u_{3j}^{G}\}$ of GMRES. In general, aNGMRES($\infty$,$p$) can be used as an alternative to GMRES.
@@
Figure 5: Convergence history for Example 4.3: aNGMRES($\infty$,$3$) vs. GMRES.
@@ 5 Conclusion
In this work, we analyze one step NGMRES($m$), where the $m+1$ iterates are obtained by fixed-point iteration, then one iteration of NGMRES($m$) is applied to these iterate. Based on this, we propose alternating NGMRES with depth $m$ and periodicity $p$, denoted as aNGMRES($m,p$). The proposed method avoids to solve the least-squares problem at each iteration, which can save the computational works. We establish the relationship between aNGMRES($m,p$), full GMRES, and restarted GMRES. aNGMRES($\infty,p$) can be regarded as an alternative to GMRES for solving linear systems. For finite $m$, the iterates of aNGMRES($m,m+1$) and restarted GMRES (GMRES($m+1$)) are identical at the iteration index $jp$. These findings increases our understanding of NGMRES. Finally, we present a convergence analysis of aNGMRES. Numerical experiments validate our theoretical findings.

### v2

@@ 4 Numerical results
In this section, we begin by discussing practical aspects of using aNGMRES, including the selection of the window size $m$, possible strategies for solving the associated least-squares problems, and the choices of norms used in those problems. We then present some examples to confirm our theoretical findings and further explore the convergence behavior of aNGMRES in cases where GMRES stagnates. In our numerical test, we use MATLAB’s backslash operator as the solver for the least-squares problems. Finally, we demonstrate the effectiveness of aNGMRES in solving a challenging nonlinear PDE boundary-value problem. A comparison with Anderson acceleration is provided.
4.1 Practical considerations
In practice, the choice of $m$ in Algorithm 1 is often problem-dependent; however, it should remain modest in size—generally within a few tens, say between 1 and 20—to ensure efficiency and stability. A large value of $m$ increases the memory requirements for solving the associated least-squares problem (LSP) and may lead to ill-conditioned systems. We note that alternative norms, such as the $\ell^{1}$ or $\ell^{\infty}$ norms, can be used in the optimization problem (5). In this work, we focus on the $\ell^{2}$ norm, under which the minimization problem can be reformulated as a linear least-squares problem and solved efficiently. A thorough examination of the parameter $m$ and the impact of different norm choices is reserved for future research.
Choosing an appropriate solution method for the LSP, (5), is a key consideration in implementing NGMRES algorithm. Although NGMRES was originally proposed in 1997, to the best of our knowledge, its practical implementation has not been thoroughly discussed in the literature. As the iterates $u_{k}$ approach the exact solution $u^{*}$, the matrix $\mathcal{D}_{k}$ in (9) may become ill-conditioned. In this case, solving the normal equations corresponding to (9) or using (10) is undesirable, as this approach squares the condition number and can lead to numerical instability.
Notably, Anderson acceleration (AA) [1] also involves solving an LSP and the issue of ill-conditioning has been extensively addressed in that context. For example, Fang and Saad [9] suggest using rank-revealing QR decomposition or a truncated singular value decomposition. Another strategy involves QR factor-updating techniques combined with adaptive control of the window size parameter $m$ to drop columns if the coefficient matrix of the LSP becomes too ill-conditioned. Moreover, recent work has proposed new approaches to improve the efficiency of AA. For instance, Lockhart et al. [13] introduced three low-synchronization orthogonalization algorithms within AA, which reduce the number of global reductions per iteration to a constant of 2 or 3, independent of the parameter $m$. Parallel implementations of AA have been investigated [14]. These techniques appear well-suited for adaptation within the NGMRES framework and may facilitate more robust and efficient implementations, which we intend to explore in future research.
4.2 Validation
@@
To confirm Theorem 3.1, we take $m=3$ and $p=m+1=4$. We report the convergence history of aNGMRES(3,4) and GMRES(4). From Theorem 3.1, we know that the $(4j)$th iterate of the two methods is the same, which is confirmed in Figure 1.
Figure 1: Residual norm convergence history of aNGMRES(3,4) vs. GMRES(4) for Example 4.1.
@@
Figure 2: Residual norm convergence history for Example 4.1. Left: aNGMRES($\infty$,4) vs. GMRES. Right: aNGMRES($\infty$,5) vs. GMRES.
@@
Consider $\mathcal{A}x=b$ [15], where $\mathcal{A}$ is block diagonal with $q$ blocks. Each block has the same structure as $A$ defined in (31). The $k$th block has dimension $\ell k$, where $\ell\in\mathbb{N}$.
Let us consider $\ell=3$ and $q=5$. Then, $\mathcal{A}\in\mathbb{R}^{45\times 45}(3+6+9+12+15=45)$. Note that $\mathcal{A}$ is invertible. We consider a zero initial guess, and construct $b$ by setting to one all the entries whose index coincides with the row index of matrix $\mathcal{A}$ where a new block starts, i.e., the nonzero components of vector $b$ are at indexes of 1, 4, 10, 19, 31. $\nu(A,r_{0})=30$. In this example, the full GMRES will stagnate for $\ell$ consecutive iterations.
@@
Figure 3: Residual norm convergence history for Example 4.2: aNGMES(2,3) vs. GMRES(3).
In Figure 4, we consider the performance of full GMRES and aNGMRES($\infty$,$p$) with $p=1,2,3,4$. Note that aNGMRES($\infty$,1) is NGMRES($\infty$). It can be seen that GMRES stagnates for three consecutive iterations, while aNGMRES($\infty$,1) stagnates forever, i.e., $u_{k}=u_{0}$ for all $k$. For aNGMRES($\infty$,2), we see the norm of the residual sequence forms an alternating periodic sequence with a period of 2. However, we see that $\|r_{3j}\|$ of aNGMRES($\infty$,3) is the same as $\|r_{3j}^{G}\|$ (i.e., $x_{3j}=x_{3j}^{G}$), and aNGMRES($\infty$,3) converges to the exact solution at the 30th iteration, which is true for GMRES. Moreover, aNGMRES($\infty$, 4) converges at the 40th iteration.
@@
Figure 4: Residual norm convergence history for Example 4.2: aNGMRES($\infty$,$p$) with $p=1,2,3,4$ vs. GMRES.
@@
We consider a 2D Laplacian with finite difference scheme and construct the coefficient matrix by using the MATLAB command $A=delsq(numgrid(^{\prime}S^{\prime},n+2))$ with $n=64$. We take $Au=b$ with $b$ all the ones and a random initial guess.
@@
In this example, the fixed-point iteration $q(u_{k+1})=Mu_{k}+b$ with M=IAM=I-A does not converge. To validate Theorem 3.2, we compare aNGMRES($\infty$,33) with GMRES in Figure 5. We see that the subsequence $\{u_{3j}\}$ of the aNGMRES($\infty$,33) method coincides with the sequence $\{u_{3j}^{G}\}$ of GMRES. In general, aNGMRES($\infty$,$p$) can be used as an alternative to GMRES.
@@
Figure 5: Residual norm convergence history for Example 4.3: aNGMRES($\infty$,33) vs. GMRES.
4.3 Applications
Although this work primarily emphasizes the theoretical and practical aspects of aNGMRES in conjunction with GMRES, we now present a numerical study demonstrating the application of aNGMRES to a nonlinear PDE boundary-value problem.
Example 4.4.
We consider the following 2D Bratu problem [24]
$\Delta u+\lambda e^{u}=0,$ $\quad\text{in}\quad\Omega=(0,1)\times(0,1),$
$u=0,$ $\quad\text{on}\quad\partial\Omega.$
In this experiment, we used a centered-difference discretization on an $N\times N$ grid ($h=1/(N+1)$). This results in a system of the form
$Au=\lambda e^{u}:=f(u),$ (32)
where $A$ is the discrete negative Laplacian matrix. We define g(u)=Auf(u)g(u)=Au-f(u), and look for solution of $g(u)=0$. The fixed-point iteration that we consider is
$q(u_{k})=u_{k}-P^{-1}(Au_{k}-f(u_{k})).$ (33)
where $P=I$ or $P={\rm diag}(A)$. We take $\lambda=6$ in the Bratu problem and use the zero initial approximation.
Our experiments were conducted in MATLAB. Since MATLAB timings may not accurately reflect performance in other computational environments and are not fully optimized, we have chosen to focus on reporting iteration counts rather than timing. Implementing the alternating NGMRES method in more advanced environments, such as C or Python, and comparing its performance with other approaches represents a valuable direction for future research. We perform a number of numerical tests by varying the window size $m$, periodicity $p$, and mesh size $h$. The stopping criterion is $\|r_{k}\|\leq 10^{-8}$.
When $P=I$, the fixed-point iteration defined by (33) corresponds to the classical Picard iteration and it diverges in this case. In our experiments, both AA and aNGMRES also exhibit a divergent behavior under this setting. Interestingly, NGMRES($m$) converges for certain values of $m$, such as 2, 6, 10, 20. In Figure 6, we report the performance of NGMRES(2), NGMRES(10), NGMRES(15), and NGMRES(20). Among these, NGMRES(15) demonstrates the best performance, followed by NGMRES(10) and NGMRES(2). Surprisingly, NGMRES(20) performs nearly the same as NGMRES(2). It may suggest that excessively large window sizes $m$ may not yield additional benefits and could even hinder convergence efficiency.
Figure 6: Residual norm convergence history for Example 4.4 for NGMRES($m$) with $h=1/129$ and $P=I$, and $m=2,10,15,20$.
When $P={\rm diag}(A)$, the fixed-point iteration defined by (33) exhibits very slow convergence. To investigate this behavior, we focus on this iteration and perform a series of tests using AA($m$), NGMRES($m$), aAA($m,p$) and aNGMRES($m,p$), systematically varying the parameters $m$ and $p$.
In Figure 7, we report the residual norm convergence history of AA($m$), NGMRES($m$), aAA($m,p$) and aNGMRES($m,p$) for $(m,p)=(2,2)$ and $(m,p)=(5,2)$. The left panel of Figure 7 shows that NGMRES(2) and aNGMRES(2,2) outperform both AA(2) and aAA(2,2), and AA and aAA stagnate. Among them, NGMRES(2) requires fewer iterations than aNGMRES(2,2) to reach the stopping criterion.
In contrast, the right pane of Figure 7, corresponding to $(m,p)=(5,2)$, shows that NGMRES(5) and aNGMRES(5,2) exhibit nearly identical convergence behavior, both outperforming AA(5) and aAA(5,2). In this case, we expect aNGMRES to achieve a lower runtime than NGMRES. Additionally, we see aAA(5,2) converges much faster than AA(2).
In Figure 8, we consider $(m,p)=(15,2)$ and $(m,p=(15,5)$. From the left panel of Figure 8, we observe that NGMRES(15) and aNGMRES(15,2) exhibit a similar convergence behavior and outperform both AA(15) and aAA(15,2). In the right panel of Figure 8, NGMRES(15), aNGMRES(15,5) and aAA(15,5) show comparable convergence histories, all outperforming AA(15).
In Figure 9, we examine the performance of the methods with parameter settings $(m,p)=(20,2)$ and $(m,p=(20,5)$. Notably, AA(20) performs worse than AA(10), as shown in Figure 8. GMRES(20), aNGMRES(20,2), and aAA(20,2) exhibit similar performance. However, aAA(20,5) shows a slight improvement over both GMRES(20), aNGMRES(20,5). NGMRES(20) and NGMRES(15) demonstrate comparable performance, as illustrated in Figure 8.
Figure 7: Residual norm convergence history for Example 4.4 for AA($m$), NGMRES($m$), aAA($m,p$) and aNGMRES($m,p$) with $h=1/129$ and $P={\rm diag}(A)$. Left: $(m,p)=(2,2)$. Right: $(m,p)=(5,2)$.
Figure 8: Residual norm convergence history for Example 4.4 for AA($m$), NGMRES($m$), aAA($m,p$) and aNGMRES($m,p$) with $h=1/129$ and $P={\rm diag}(A)$. Left: $(m,p)=(15,2)$. Right: $(m,p)=(15,5)$.
Figure 9: Residual norm convergence history for Example 4.4 for AA($m$), NGMRES($m$), aAA($m,p$) and aNGMRES($m,p$) with $h=1/129$ and $P={\rm diag}(A)$. Left: $(m,p)=(20,2)$. Right: $(m,p)=(20,5)$.
In Figure 10, we examine the performance of aAA(20,2) and aNGMRES(20,2) by varying the mesh size $h=1/(N+1)$. The tests are labeled as aAA-N and aNGMRES-N, corresponding to $N=32,64,128,256$. The number of iterations required to meet the stopping criterion for aAA($20,2$) are 114, 261, 486, and 1268, while for aNGMRES($20,2$) they are 80, 196, 470, and 1074, respectively. The computational cost per iteration does not make a big difference between the two methods. We also report the iteration counts for aAA($20,5$) and aNGMRES($20,5$). The iteration counts for aAA($20,5$) are 70, 175, 405, and 1135, and for aNGMRES($20,5$) are 75, 195, 475, 1085.
Figure 10: Residual norm convergence history for Example 4.4 for aAA and aNGMRES with $P={\rm diag}(A)$, varying NN. Left: aAA($20,2$) and aNGMRES($20,2$). Right: aAA($20,5$) and aNGMRES($20,5$).
In summary, this example demonstrates that NGMRES can converge even for noncontractive fixed-point iterations. The adaptive Anderson aAA significantly outperforms standard AA. With appropriate choices of $m$ and $p$, NGMRES($m$), aAA($m,p$), NGMRES($m$), and aNGMRES($m,p)$ exhibit comparable performance. However, for large values of $m$, both AA($m$) and NGMRES($m$) ($m\geq 15$) show limited improvement.
This is a preliminary case study. In future work, we will explore the application of aNGMRES to more complex systems, particularly nonlinear problems that pose significant computational challenges. A detailed comparison with Anderson Acceleration will be conducted to assess its performance relative to NGMRES or aNGMRES.
@@ 5 Conclusion
In this work, we analyze one-step NGMRES($m$) applied to linear systems, where the $m$ iterates are obtained by Richardson iteration, then one iteration of NGMRES($m$) is applied to these iterates. Based on this, we propose alternating NGMRES with depth $m$ and periodicity $p$, denoted as aNGMRES($m,p$). The proposed method avoids solving the least-squares problem at each iteration, which can save computational work. We establish the relationship between aNGMRES($m,p$), full GMRES, and restarted GMRES. aNGMRES($\infty,p$) can be regarded as an alternative to GMRES for solving linear systems. For finite $m$, the iterates of aNGMRES($m,m+1$) and restarted GMRES (GMRES($m+1$)) are identical at the iteration index $jp$. These findings increase our understanding of NGMRES. Finally, we present a convergence analysis of aNGMRES applied to linear systems, using Richardson iteration as the underlying fixed-point scheme. Numerical experiments validate our theoretical findings.